OpenStudy (anonymous):

Find the 11th partial sum of the summation of negative 7i plus 22, from i equals 1 to infinity.

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty}-7i+22\]

zepdrix (zepdrix):

\[\Large\rm \sum_{i=1}^{11}-7i+22=\sum_{i=1}^{11}-7i+\sum_{i=1}^{11}22\] \[\Large\rm =\sum_{i=1}^{11}-7i+11(22)\]

zepdrix (zepdrix):

That second pard simplifies down nicely. For the first part, \[\Large\rm -7\sum_{i=1}^{11}i=-7(1+2+...+11)\]

OpenStudy (anonymous):

mk :) then what

zepdrix (zepdrix):

\[\Large\rm =-7(1+2+...+11)+11(22)\]Just bust our your calculator from that point I guess :

OpenStudy (anonymous):

what are the ... for?

zepdrix (zepdrix):

To show a continuation. I didn't want to type out 1+2+3+4+5+6+7+8+9+10+11. That's what the dots are implying.

OpenStudy (anonymous):

ohhhh ok thanks

OpenStudy (kaylaprincess):

i dont get what the answer is?